OpenStudy (anonymous):

There are 15 children playing. 6 of the children have brown hair. There are 5 more children with black hair than red hair. How many children have red hair?

OpenStudy (anonymous):

15-6=9 2+5=7 Two red heads

OpenStudy (anonymous):

Tks

OpenStudy (anonymous):

This does not add up to 15 children....

OpenStudy (anonymous):

7+2+6=15